Hi. I'm needing help in Algebra.
Here's my homework problem:
List (a) the coordinates of any y-intercept and (b) the coordinates of any x-intercept. Do not graph.
−5x + 6y = 100
Thank you!

Answer by nerdybill(6958)

List (a) the coordinates of any y-intercept and (b) the coordinates of any x-intercept. Do not graph.
−5x + 6y = 100
.
set x=0 to find y-intercept:
−5x + 6y = 100
−5(0) + 6y = 100
0 + 6y = 100
y = 100/6
y = 50/3
y-intercept is at (0, 50/3)
.
set y=0 to find x-intercept:
−5x + 6y = 100
−5x + 6(0) = 100
−5x + 0 = 100
x = -20
x-intercept is at (-20, 0)